OpenStudy (anonymous):

Find the arc length of the curve r(t)=<6t^2, 2sqrt(6)t, ln t>, 1

OpenStudy (zarkon):

\[\int\limits_{a}^{b}\sqrt{\left(\frac{dx(t)}{dt}\right)^2+\left(\frac{dy(t)}{dt}\right)^2+\left(\frac{dz(t)}{dt}\right)^2}dt\]